How much of a 5% acid solution and a 20% acid solution must be mixed inorder to obtain 3 gallons of a 10% acid solution? I got the answer 2 gallons of 5% and 1 gallon of 10% but I don't know why.
Please show me step by step to get the answer. Thanks

Answer by jorel1380(3719)

3 gals. of 10% solution=3x.1 or .3 gals pure acid.
-----------------------
x+y=3
.20x+.05y=.3
20x+5y=30
5x+5y=15
15x=15
x=1
y=2
You need 1 gal. of 20% solution, and 2 gals. of 5% solution for the mix..